### v2

# FlowMalTrans: Unsupervised Binary Code Translation for Malware Detection Using Flow-Adapter Architecture

Abstract
Applying deep learning to malware detection has drawn great attention due to its notable performance. As malware spreads across a wide range of Instruction Set Architectures (ISAs), it is important to extend malware detection capacity to multiple ISAs. However, training a deep learning-based malware detection model usually requires a large number of labeled malware samples. The process of collecting and labeling sufficient malware samples to build datasets for each ISA is labor-intensive and time-consuming. To reduce the burden of data collection, we propose to leverage the ideas of Neural Machine Translation (NMT) and Normalizing Flows (NFs) for malware detection. Specifically, when dealing with malware in a certain ISA, we translate it to an ISA with sufficient malware samples (like X86-64). This allows us to apply a model trained on one ISA to analyze malware from another ISA. We have implemented and evaluated the model on seven ISAs. The results demonstrate the high translation capability of our model, enabling superior malware detection across ISAs. We have made our code and dataset open-sourced available [^1] https://github.com/mhu16419/FlowMalTrans.
FlowMalTrans: Unsupervised Binary Code Translation for Malware Detection Using Flow-Adapter Architecture
Minghao Hu George Mason University mhu20@gmu.edu          Junzhe Wang George Mason University jwang69@gmu.edu          Weisen Zhao George Mason University wzhao9@gmu.edu
Qiang Zeng George Mason University zeng@gmu.edu          Lannan Luo 🖂 George Mason University lluo4@gmu.edu
1 Introduction
Malware is defined as software intended to damage computers or associated systems Preda et al. (2008). In recent years, many malware detection tools have been developed Xie et al. (2024), and their success largely depends on the underlying techniques. Signature-based detection identifies malware by matching patterns from known malware families Sathyanarayan et al. (2008), but it often fails to detect altered or novel malware. In contrast, behavior-based detection analyzes program execution to identify suspicious behavior Liu et al. (2011), yet this method lacks scalability.
The use of deep learning methods like LSTM and CNNs in malware detection has garnered significant interest because of their strong performance Sewak et al. (2018); Gopinath and Sethuraman (2023). However, there are two major challenges that hinder the broader adoption of deep learning in this domain.
Challenges. Malware infects a broad range of Instruction Set Architectures (ISAs). With the rise in cyberattacks on IoT devices Chen et al. (2025); Ma et al. (2023); Tsai et al. (2024), malware has begun targeting an increasing number of ISAs Davanian and Faloutsos (2022); Caviglione et al. (2020). Hence, it is essential to enable malware detection across ISAs. Currently, there are numerous ISAs available Wang et al. (2023a). Due to this wide variety, training one model for each ISA demands substantial time and effort. In addition, deep learning models rely heavily on having enough malware samples during training, which is problematic for low-resource ISAs where such data is limited.
Our Approach. Malware is generally closed-source whose original source code is typically inaccessible. What is available is the binary form of the malware. Once disassembled, this binary can be represented in assembly language. Based on this observation, we propose leveraging Neural Machine Translation (NMT) techniques Artetxe et al. (2018) to facilitate malware analysis.
When handling a binary in a given ISA (referred to as the source ISA), we translate it to an ISA with rich malware samples, such as X86-64, which we refer to as the target ISA. Once translated, we use a model trained on the target ISA to test the translated code. This approach facilitates malware detection across multiple ISAs using a model trained exclusively on the target ISA, eliminating the need for extensive malware samples in other ISAs.
In order to capture ISA-specific syntax and semantic information, we leverage normalizing flows (NFs), a class of invertible transformations that models complex distributions while maintaining tractability. Traditional translation models often struggle with capturing complex dependencies and handling the variation in data distributions. NFs address this challenge by modeling complex, high-dimensional distributions through invertible transformations, enabling more efficient learning of intricate relationships between source and target data. By doing so, we ensure a smooth and structured mapping between ISAs, leading to more precise and reliable translations. Furthermore, normalizing flows enable efficient sampling and density estimation, which improves the adaptability of our model to unseen binaries and enhances generalization across different architectures.
We design FlowMalTrans, which is a flow-adapter-based unsupervised binary code translation architecture. This design allows for robust feature alignment between different ISAs, improving translation fidelity and ultimately enhancing malware detection across ISAs. To bridge the gap between the latent space of the source and target ISA, FlowMalTrans learns an invertible transformation of binary representations from one ISA to another, facilitating robust feature alignment across different ISAs and enhancing translation fidelity. FlowMalTrans operates in a completely unsupervised manner, eliminating the need for parallel datasets. Notably, the training of FlowMalTrans does not require any malware samples and relies only on binaries compiled from the abundance of opensource programs. Despite never encountering any malware samples during training, FlowMalTrans is still capable of translating malware across ISAs with high translation quality.
Results. We have implemented FlowMalTrans, and evaluated its performance on seven ISAs, including X86-64, i386, ARM32, ARM64, MIPS32, PPC32 and M68K. FlowMalTrans achieves higher BLEU score than the baseline method UnsuperBinTrans Ahmad and Luo (2023). Furthermore, we apply FlowMalTrans to the malware detection task and achieve exceptional results, highlighting its superior translation quality, which enables highly effective malware detection across ISAs. Below we highlight our contributions:
  • We propose FlowMalTrans, a novel unsupervised approach to translate binaries across different ISAs. FlowMalTrans addresses the data scarcity issue by enabling the detection of malware in low-resource ISAs using a model trained on the high-resource ISA. Importantly, the training of FlowMalTrans does not require any malware samples, yet it is still capable of translating malware across ISAs and achieves high translation quality.
  • We leverage normalizing flows (NFs) to capture ISA-specific features. With NFs, FlowMalTrans captures code semantics and and properly decodes instruction sequences.
  • The training of FlowMalTrans does not require any malware samples, yet the model is still capable of translating malware across ISAs and achieves high translation quality.
  • We expand the evaluation to cover a broader range of ISAs compared to prior work, showcasing FlowMalTrans’s wider applicability and improved performance.
2 Related Work
2.1 Malware Detection
Signature-Based Detection. Conventional malware detection methods primarily use signature-based approaches, working by recognizing malicious code patterns Sebastio et al. (2020); Rohith and Kaur (2021); Sathyanarayan et al. (2008); Behal et al. (2010). However, these methods are ineffective against newly developed or altered malware that modifies its code to avoid being identified.
Behavior-Based Detection. This approach detects malware by examining program behavior during execution Liu et al. (2011); Burguera et al. (2011); Aslan et al. (2021); Saracino et al. (2016). However, it may produce false negatives when malicious actions are not activated during observation.
Deep Learning-Based Detection. Deep learning has been widely used for detecting malware. A range of models, including CNNs and LSTMs, have been utilized Sewak et al. (2018); Gopinath and Sethuraman (2023); Wang et al. (2024); He et al. (2023); Lei et al. (2022); Zhang et al. (2024); Zhao et al. (2025). Nonetheless, these models generally demand substantial datasets of malware samples for training. Consequently, most existing methods concentrate on widely used ISAs like X86 and ARM, leaving low-resource ISAs relatively unaddressed.
2.2 Normalizing Flows
Normalizing flows (NFs) are a family of generative models that transform a simple base distribution into a complex target distribution via a series of invertible transformations Ho et al. (2019). The key idea is to apply a sequence of differentiable mappings to a latent variable. Different types of NFs include coupling layers Dinh et al. (2014), autoregressive flows Papamakarios et al. (2017), and continuous normalizing flows Chen et al. (2018).
The flow adapter architecture has been explored in neural machine translation (NMT), such as variational NMT Calixto et al. (2019); Ma et al. (2019); Eikema and Aziz (2019). Flow adapters leverage NFs to model complex posterior distributions in sequence-to-sequence models, allowing for more flexible latent representations and enhancing the robustness of translation Shu et al. (2020).
2.3 Source Code Translation
Source code translation converts source code in one language to another. Early work uses transpilers or transcompilers Andrés and Pérez (2017); Tripuramallu et al. (2024), which rely on handcrafted rules. However, they produce unidiomatic translations that prove hard for human programmers to read. Another issue is incomplete feature support, resulting in improper translations.
Deep learning have introduced new approaches to source code translation Roziere et al. (2020); Weisz et al. (2021); Lachaux et al. (2020). However, since malware is closed-source, source code translation is inapplicable for malware.
2.4 Program Analysis-Based Binary Code Translation
Several approaches leverage program analysis techniques to convert binary code from one ISA to another. They can be broadly categorized into static analysis-based and dynamic analysis-based. Static analysis-based translation analyzes and translates the binary code before execution Shen et al. (2012); Cifuentes and Van Emmerik (2000). But ensuring all paths are accurately translated is challenging.
Dynamic analysis-based translation performs translation during execution Chernoff et al. (1998); Ebcioglu et al. (2001), but it requires sophisticated runtime environments. Plus, both approaches face challenges related to architecture-specific features and encounter difficulties in achieving accuracy, performance, and compatibility.
Summary. We propose an unsupervised binary code translation approach leveraging deep learning techniques. The state of the art, UnsuperBinTrans Ahmad and Luo (2023), is the first and only existing work applying deep learning to binary code translation. However, it has several limitations, which our model FlowMalTrans aims to overcome (Section 3.2). Our evaluation demonstrates that FlowMalTrans outperforms UnsuperBinTrans in binary code translation and achieves superior malware detection performance.
3 Overview
Figure 1: Applying FlowMalTrans to detect malware in a source ISA using a malware detection model trained on the target ISA (e.g., X86-64).
3.1 Motivation
Malware propagates over various instruction set architectures (ISAs), highlighting the importance of enabling malware detection across multiple ISAs. Nonetheless, training deep learning models generally depends on extensive datasets of malware samples. This requirement becomes more difficult to fulfill with the diversity of ISAs, as low-resource ISAs frequently lack adequate malware samples.
To alleviate the challenges of data collection and tackle the problem of limited data, we propose translating binaries from a low-resource ISA to a high-resource ISA. Thus malware detection model trained on the high-resource ISA can be applied to the translated binaries, helping to overcome the data scarcity faced by low-resource ISAs.
3.2 Limitations of the State of the Art
UnsuperBinTrans Ahmad and Luo (2023) is the first and only existing work that applies deep learning techniques to binary code translation. While InnerEye Zuo et al. (2018) applies NMT techniques for binary code similarity comparison, it does not perform binary code translation across ISAs. Instead, it employs two encoders from NMT models to generate embeddings for binary pairs. UnsuperBinTrans, however, has several limitations, which our model aims to overcome.
Shared Encoder vs. Separated Encoders. UnsuperBinTrans employs a shared encoder for multiple ISAs, which limits the model’s ability to capture ISA-specific syntax and semantics, leading to suboptimal translation. In contrast, FlowMalTrans employs a separate encoder for each ISA, allowing each to learn specialized representations.
RNN vs. Transformer. UnsuperBinTrans relies on RNNs, which suffer from vanishing gradients and difficulties in capturing long-range dependencies. In contrast, FlowMalTrans leverages Transformer, which employ self-attention mechanisms to model long-range dependencies.
ISA-Agnostic vs. ISA-Specific Latent Representations. UnsuperBinTrans learns latent representations for binaries that are ISA-agnostic by abstracting away ISA-specific details. In contrast, FlowMalTrans learns ISA-specific latent representations that preserve crucial architectural distinctions while maintaining alignment with a shared cross-ISA space. This is achieved by employing NFs. By transforming the source-ISA representations into the target-ISA representations via NFs, the decoder can leverage the representations to generate better aligned target-ISA binary code.
Two ISAs vs. Broader ISAs. UnsuperBinTrans is limited to only two ISAs (X86-64 and ARM32), whereas FlowMalTrans extends its coverage, supporting seven ISAs.
3.3 Model Applications
Figure 1 shows the workflow of applying FlowMalTrans to detect malware across ISAs. In step 1, we train FlowMalTrans to translate binaries from the source ISA to the target ISA (such as X86-64). The training dataset can be constructed using various available opensource programs. The training of FlowMalTrans does not require any malware samples. It should be noted that malware is typically a closed-source program; thus cross-compilation that generates binaries across ISAs from source code does not apply to malware.
In step 2, we train a deep learning model using a task-specific dataset containing malware and benign samples in the target ISA.
Finally, in step 3, when dealing with a binary in the source ISA, we use FlowMalTrans to translate the binary to the target ISA and reuse the malware detection model trained on the target ISA to test the translated code for detecting malware.
Note that due to the scarcity of malware samples in a low-resource ISA, directly training a robust malware detection model on such an ISA is challenging. Our approach overcomes this limitation through code translation, making robust malware detection feasible for low-resource ISAs.
4 Model Design
4.1 Instruction Normalization
A binary, after being disassembled, is represented as a sequence of instructions. An instruction includes an opcode and zero or more operands. We regard opcodes/operands as words and basic blocks as sentences. A basic block is a straight-line sequence of instructions with no branches inside it.
In Natural Language Processing (NLP), the out-of-vocabulary (OOV) issue is a well-known problem. To mitigate the OOV problem, we employ the normalization strategy.
  • (R1): We use IDA Pro IDA (2023) to disassemble binaries, which generates dummy names Dummy name (2023). We replace dummy names with their respective prefixes. For example, off_ and seg_ represent offset pointer value and segment address value. They are replaced with <OFF> and <SEG>.
  • (R2): Function names are replaced by <FUNC>.
  • (R3): Number and hexadecimal constants are replaced by <VALUE> and <HEX>, respectively.
We provide examples in Appendix A.2 that illustrate how these normalization rules are applied to assembly code across different ISAs.
Figure 2: Model architecture of FlowMalTrans, which contains a pair of flow adapters for modeling the distributions of basic block representations in the source ISA and target ISA, respectively.
4.2 Normalizing Flows
We consider basic blocks as sentences, and instructions as words. The prior state-of-the-art, UnsuperBinTrans, assumes that the latent representations of source and target basic blocks share a common semantic space, and learns ISA-agnostic latent representations, which abstract away ISA-specific details. However, this can be overly restrictive, as different ISAs exhibit distinct syntactic and structural characteristics that are not always easily aligned in a shared latent space. By enforcing ISA-agnostic representations, UnsuperBinTrans may lose critical ISA-specific features necessary for accurate translation, leading to suboptimal performance.
Our model FlowMalTrans assumes the representations are different due to ISA-specific characteristics. Thus, they are modeled separately for each ISA, which makes it possible to better capture basic-block semantics in an ISA-specific manner. Specifically, FlowMalTrans uses a pair of NFs (i.e., source flow and target flow) to model the distributions of basic-block latent representations in the source and target ISA. During translation, a latent code transformation is performed to transform the source representation $z_{src}$ into the target representation $z_{tgt}$, which is fed to the decoder to generate a better basic block in the target ISA. Figure 2 shows how NFs are employed in our model.
Modeling Latent Representations by NFs. We use a pair of NFs to model the distribution of the basic block latent representations in different ISAs, i.e., $p_{z_{src}}(z_{src})$ and $p_{z_{tgt}}(z_{tgt})$, where $z_{src}$ and $z_{tgt}$ are the latent representation of the source basic block and target basic block, respectively. Through NFs, we transform the distributions of the source and target representations to the base distribution $\epsilon$ (e.g., standard normal distribution), which can be viewed as the “true” underlying semantic space, abstracting away from ISA specifics. We denote such mappings as $\textbf{G}_{(z_{src}\rightarrow\epsilon)}$ and $\textbf{G}_{(z_{tgt}\rightarrow\epsilon)}$.
Designing NFs. We use Real-valued Non-volume Preserving (RealNVP), a type of NFs designed for efficient density estimation and sampling Dinh et al. (2022), to build the source and target flow.
Two key components of RealNVP are: Multi-scale Architecture and Affine Coupling Layer. With them, RealNVP can: (1) capture ISA-specific representations for each ISA at both local and global scales; (2) transform these encoded latent representations from one ISA to another invertibly; and (3) train using maximum likelihood estimation (MLE) with a tractable Jacobian determinant.
We construct FlowMalTrans with two RealNVP models for the source flow and target flow. Each RealNVP model consists of three sequential flows to perform the latent code transformation. We explicitly model the source and target blocks with $K$ sequential flows (where $K=3$ in our case):
$p_{z_{src}}(z_{src})=p_{\varepsilon}(\varepsilon)\prod_{i=1}^{K}\left|\det\frac{\partial f_{src}^{(i)}(z^{(i)})}{\partial z^{(i)}}\right|^{-1}$ (1)
$p_{z_{tgt}}(z_{tgt})=p_{\varepsilon}(\varepsilon)\prod_{i=1}^{K}\left|\det\frac{\partial f_{tgt}^{(i)}(z^{(i)})}{\partial z^{(i)}}\right|^{-1}$ (2)
where $p_{\varepsilon}(\varepsilon)$ is a base distribution. We select standard Gaussian distribution $\mathcal{N}(0,1)$ as our base distribution here. $f_{src}^{(i)}$ is the $i$-th transformations for the source blocks. $z^{(i)}$ is the intermediate variable where $z^{(1)}=\varepsilon$ and $z^{(K)}=z_{src}$.
We denote the mapping process in Equation (1) as $\textbf{G}_{(\varepsilon\rightarrow z_{src})}$ and Equation (2) as $\textbf{G}_{(\varepsilon\rightarrow z_{tgt})}$. Given the invertibility feature of NFs, the mappings are also invertible. Thus, we have: $\textbf{G}_{(\varepsilon\rightarrow z_{src})}=\textbf{G}^{-1}_{(z_{src}\rightarrow\varepsilon)}$ and $\textbf{G}_{(\varepsilon\rightarrow z_{tgt})}=\textbf{G}^{-1}_{(z_{tgt}\rightarrow\varepsilon)}$. To achieve latent code transformation from source to target, we can formalize the transformation process as the composition of $\textbf{G}_{(z_{src}\rightarrow\varepsilon)}$ and $\textbf{G}_{(\varepsilon\rightarrow z_{tgt})}$:
$\textbf{G}_{(z_{src}\rightarrow z_{tgt})}=\textbf{G}_{(z_{src}\rightarrow\varepsilon)}\circ\textbf{G}_{(\varepsilon\rightarrow z_{tgt})}$ (3)
Therefore, we can learn a transformation $\textbf{G}_{(z_{src}\rightarrow z_{tgt})}$ by learning the transformation $\textbf{G}_{(z_{src}\rightarrow\varepsilon)}$ and $\textbf{G}_{(\varepsilon\rightarrow z_{tgt})}$. We also notice that $\textbf{G}_{(z_{src}\rightarrow z_{tgt})}$ and $\textbf{G}_{(z_{tgt}\rightarrow z_{src})}$ are invertible because they are compositions of two invertible mappings. Thus, We can follow a similar procedure and learn $\textbf{G}_{(z_{tgt}\rightarrow z_{src})}$.
4.3 Encoder & Decoder for Code Translation
For binary code translation, different ISAs exhibit significant syntactic and semantic differences. To capture and align these variations, we propose a separated-encoder design, where each ISA is assigned a dedicated encoder. The source encoder and decoder are responsible for encoding and generating basic blocks in the source ISA, while the target encoder and decoder handle basic blocks in the target ISA. FlowMalTrans employs a multi-layer bi-directional Transformer Vaswani (2017) to construct the encoders and decoders for the source and target ISA, as shown in Figure 2.
Encoding. Here we describe how the source encoder encodes the source basic block and generates the source basic block representation $z_{src}$. A similar procedure is applied to the target basic block representation generation.
The source encoder processes the source basic block $\textbf{B}_{src}=\{\boldsymbol{b}_{0},\cdots,\boldsymbol{b}_{S}\}$, and generates the hidden representations $\{\boldsymbol{h}_{0},\cdots\boldsymbol{h}_{S}\}$. The basic block-level representation $z_{src}$ is computed using Equation (4). The target encoder follows a similar process to encode the target basic block.
$z_{src}=\boldsymbol{W}($ $\ \text{max-pool}([\boldsymbol{h}_{0},\dots,\boldsymbol{h}_{S}])$
$+\text{mean-pool}([\boldsymbol{h}_{0},\dots,\boldsymbol{h}_{S}])$
$+\boldsymbol{h}_{0})$ (4)
Cross-lingual Translation. To enable the decoder to better leverage ISA-specific latent representations, we perform a latent code transformation. For example, as shown in Figure 2, after calculating $z_{src}$ using Equation (4), we employ the source flow to transform $z_{src}$ to $\epsilon$, and then the target flow to transform $\epsilon$ to $z_{tgt}$, which is in the target latent representation space. Then, the target decoder can decode $z_{tgt}$ to generate the target basic block.
In contrast, the prior work UnsuperBinTrans indiscriminately uses the same representational space for both source and target, without performing latent code transformation.
Decoding. To capture the semantics and mitigate improper alignments between the source and target ISA, we follow the procedure in Setiawan et al. (2020) to generate the decoded representation:
$\boldsymbol{o}_{i}=(1-\boldsymbol{g}_{i})\odot\boldsymbol{s}_{i}+\boldsymbol{g}_{i}\odot\boldsymbol{z}\vskip-3.0pt$ (5)
where $\boldsymbol{g}_{i}=\sigma([\boldsymbol{s}_{i};\boldsymbol{z}])$, $\sigma(\cdot)$ is the sigmoid function. $\boldsymbol{o}_{i}$ denotes Hadamard product between two tensors. The values in $\boldsymbol{g}_{i}$ control the contribution of $\boldsymbol{z}$ to $\boldsymbol{o}_{i}$.
4.4 Model Training
We first employ the causal language modeling (CLM) and masked language modeling (MLM) objectives to pretrain the encoders and decoders. We then train FlowMalTrans in an unsupervised manner using three objectives: denoising auto-encoding (DAE), back translation (BT), and maximum likelihood estimation (MLE). The DAE reconstructs a basic block from its noised version, as illustrated in steps 1 and 2 in Figure 2. The BT process, shown in steps 3 and 4, involves performing the latent code transformation twice while jointly training the encoders and decoders. For example, in BT for the source ISA, the transformation first occurs in the source-to-target direction, followed by the target-to-source direction. Appendix A.1 provides further details.
Maximum Likelihood Estimation (MLE). The NFs are directly trained by maximum likelihood estimation (MLE) of basic-block level latent representations. The objective can be formulated as:
$\mathcal{L}_{MLE}(\textbf{G}_{(z_{src}\rightarrow\varepsilon)})=E_{z\sim p_{z_{src}}}[\log p_{z_{src}}(z)]$ (6)
where $E_{z\sim p_{z_{src}}}$ is approximated by sampling mini-batches of latent representations generated by the encoder during training. By minimizing the loss, we can construct $\textbf{G}_{(z_{src}\rightarrow z_{tgt})}$ and $\textbf{G}_{(z_{tgt}\rightarrow z_{src})}$.
Trained with the DAE, BT, and MLE objectives, FlowMalTrans effectively translates basic blocks across ISAs. During testing, it translates each basic block of a given binary from the source ISA and then concatenates the translated blocks to form a translated binary for the target ISA.
5 Evaluation
5.1 Experimental Settings
We implement FlowMalTrans using Transformer with $64$ hidden units, $4$ heads, ReLU activations, a dropout rate of $0.1$, and learned position embeddings. Appendix A.4 presents the details. All the experiments were conducted on a computer with a $64$-bit 3.6 GHz Intel Core i9-CPU, a Nvidia GeForce RTX 4090 and 64GB RAM.
Model Comparison. We consider four baseline models and one Same-ISA model for comparison.
  • Baseline Model 1: UnsuperBinTrans Ahmad and Luo (2023) is the first and only existing work that focuses on binary code translation leveraging deep learning (Section 3.2).
  • Baseline Model 2: UniMap Wang et al. (2023a) also aims to resolve the data scarcity issue in malware detection. UniMap learns transfer knowledge to enable the reuse of a model across ISAs. In contrast, we focus on binary code translation to enable model reuse.
  • Baseline Model 3: CrossIns2Vec Wang et al. (2024), similar to UniMap, aims to learn transfer knowledge to enable model reuse.
  • Baseline Model 4: IR-based malware detection model. Intermediate representation (IR) can abstract away ISA differences and represent binaries in a uniform style angr (2024).
  • Same-ISA model. We consider a model trained and tested on the same ISA, without employing any translation, as the Same-SIA model. As expected, this model, if trained with sufficient data, is likely to outperform a model trained on one ISA and tested on another, representing the best-case scenario.
5.2 Training FlowMalTrans
We consider seven ISAs: X86-64, i386, ARM64, ARM32, MIPS32, PPC32 and M68K. We translate binaries from other ISAs to X86-64.
Training Datasets. We collect various opensource programs, including openssl, binutils, findutils, and libgpg-error, which are widely used in prior binary analysis works Ding et al. (2019); Li et al. (2021). We compile them for each ISA using GCC with different optimization levels (O0-O3) and disassemble the binaries using IDA Pro IDA (2023) to extract basic blocks, which are normalized and deduplicated. Finally, we create a training dataset for each ISA. The adequacy of these datasets is evaluated in Appendix A.3.
Byte Pair Encoding. We use byte pair encoding (BPE) to process the datasets. We set the BPE merge times based on the principles derived from empirical experiments and theoretical insights:
  • The vocabulary size discrepancy between the source and target ISA should $15\%$. A large vocabulary discrepancy can lead to an imbalanced learning problem, resulting in inefficiencies and overfitting Gowda and May (2020).
  • The vocabulary size of each ISA should <12k<12k, to balance capturing word semantics with computational resource constraints. A large vocabulary size can negatively impact model performance due to increased complexity and sparse token distributions Jean et al. (2014).
As shown in Table 1, we can see that the vocabulary discrepancy within each pair is small, making them well-suited for training. Note that these principles are tailored to our specific scenarios.
Table 1: Vocabulary size, BPE merge times, and joint vocabulary size for each pair of ISAs.
ISA Pair Vocab. Size Vocab. Size Merge Joint Vocab.
(src $\leftrightarrow$ tgt) (src) (tgt) Time Size
i386 $\leftrightarrow$ X86-64 $7,135$ $7,104$ $10,000$ $9,688$
ARM32 $\leftrightarrow$ X86-64 $9,416$ $9,236$ $22,000$ $17,262$
ARM64 $\leftrightarrow$ X86-64 $5,142$ $4,455$ $9,000$ $7,104$
MIPS32 $\leftrightarrow$ X86-64 $10,995$ $11,620$ $14,000$ $12,685$
PPC32 $\leftrightarrow$ X86-64 $8,691$ $9,504$ $17,000$ $13,032$
M68K $\leftrightarrow$ X86-64 $7,148$ $6,484$ $9,000$ $8,386$
[…]
5.3 Testing on Binary Code Translation
We use the Bilingual Evaluation Understudy (BLEU) Papineni et al. (2002) to evaluate the translation performance, which is often used to evaluate the quality of machine-generated translations by measuring the $n$-gram overlap between the translation and reference. We set the tokenization of SacreBLEU Post (2018) to None, apply add-one smoothing, and use the default settings
Testing Datasets. We use three packages, zlib-1.2.11, coreutils-9.0, and diffutils-3.7, to test FlowMalTrans. Note that these packages are not included in the training dataset of FlowMalTrans. We compile them on the six ISAs using GCC-11.4.0 with different optimization levels (O0-O3) and use IDA Pro to disassemble them.
Translation Results. We consider six ISAs, i386, ARM32, ARM64, MIPS32, PPC32 and M68K, as the source ISAs, and X86-64 as the target ISA. For each binary $B_{1}$ in the source ISA, there exists a semantically equivalent binary $B_{2}$ in X86-64. We use FlowMalTrans to translate $B_{1}$ into X86-64, resulting in a translated binary $B_{3}$ in X86-64. The BLEU score is computed between $B_{3}$ and $B_{2}$. We report the average BLEU score for all binaries. The results are shown in Table 2. Appendix A.11 presents some translation examples.
Table 2: BLEU scores of FlowMalTrans & baseline.
ISA Pair FlowMalTrans UnsuperBinTrans
(src $\rightarrow$ tgt) (our work) (baseline)
i386 $\rightarrow$ X86-64 $0.41$ $0.35$
ARM32 $\rightarrow$ X86-64 $0.40$ $0.28$
ARM64 $\rightarrow$ X86-64 $0.37$ $0.32$
MIPS32 $\rightarrow$ X86-64 $0.36$ $0.27$
PPC32 $\rightarrow$ X86-64 $0.31$ $0.25$
M68K $\rightarrow$ X86-64 $0.39$ $0.31$
We compare FlowMalTrans to the baseline UnsuperBinTrans. We use the open-source trained model of UnsuperBinTrans for this comparison. Note that UnsuperBinTrans  focuses solely on two ISAs (X86-64 and ARM32). To ensure a comprehensive comparison, we train UnsuperBinTrans on the same mono-architecture training datasets for the additional ISA pairs. We can see that FlowMalTrans outperforms UnsuperBinTrans across all ISA pairs and demonstrates satisfactory performance. Thus, FlowMalTrans has good translation quality and can effectively translate binaries across ISAs.
The average time to translate a basic block from one ISA to x86-64 is $10^{-4}$s, and the average number of blocks in a binary is $12,320$. Thus, translating a binary takes around $1.2$s. This demonstrates the good scalability of our approach.
Table 3: Malware detection results using LSTM. In Table (a), we compare the detection performance by translating malware using FlowMalTransUnsuperBinTrans. Then we evaluate it against UniMapCrossIns2Vecand the IR-based model. In Table (b), we give the results of the Same-ISA Model, which is trained and tested on the same ISA.
(a) FlowMalTrans vs. Four baselines.
ISA Pair FlowMalTrans UnsuperBinTrans UniMap CrossIns2Vec IR-based Model
(src $\rightarrow$ tgt) (Our Work) (Baseline 1) (Baseline 2) (Baseline 3) (Baseline 4)
i386 $\rightarrow$ X86-64 0.996 $0.723$ $0.953$ - $0.819$
ARM32 $\rightarrow$ X86-64 0.981 $0.818$ $0.953$ $0.976$ $0.815$
ARM64 $\rightarrow$ X86-64 0.965 $0.638$ $0.910$ - $0.825$
MIPS32 $\rightarrow$ X86-64 0.971 $0.725$ $0.933$ $0.964$ $0.791$
PPC32 $\rightarrow$ X86-64 0.973 $0.722$ $0.919$ $0.967$ $0.639$
M68K $\rightarrow$ X86-64 0.931 $0.689$ $0.893$ - $0.476$
(b) The Same-ISA model.
Train & Test on Same-ISA
the same ISA Model
i386 $0.998$
ARM32 $0.986$
ARM64 $0.705$
MIPS32 $0.974$
PPC32 $0.849$
M68K $0.790$
5.4 Malware Detection Task
We train a malware detection model on X86-64 (where abundant malware samples are available), and reuse the trained model for other ISAs by translating binaries using FlowMalTrans  [^2] We are aware that ARM32 is also a high-resource ISA; due to its importance, our evaluation involves it. However, whether our approach can be applied to translate other ISAs to ARM32 requires further experiments..
Malware Detection Model. We use the LSTM model HaddadPajouh et al. (2018) to detect malware, designed as two layers. The model details are provided in Appendix A.5. We first extract the token embeddings from FlowMalTrans and integrate them into the input layer of LSTM. As a result, when a binary is fed into LSTM, each input token is represented as its embedding. To further enrich the task, we also apply a CNN model, with the results presented in Appendix A.6. In the following, we focus on the results related to LSTM.
Task-Specific Training Datasets. We collect $2140$, $1740$, $1581$, $1430$, $962$, $545$, and $128$ malware samples from VirusShare.com VirusShare (2020) for X86-64, i386, ARM32, MIPS32, PPC32, M68K, and ARM64, respectively. We spend significant efforts in collecting malware, especially for PPC32, M68K, and ARM64. Notably, our approach only needs malware in a high-resource ISA to train a malware detection model, significantly reducing the data collection burden for low-resource ISAs.
As we compare the results to the Same-ISA model trained and tested on the same ISA (Section 5.1), we build task-specific training datasets for each ISA. We split the malware samples into two parts: $80\%$ are used for training and $20\%$ for testing. Each training dataset also includes an equal number of benign samples, randomly selected from openssl, binutils, findutils, and libgpg-error.
We also evaluate the performance when testing on all malware samples (without comparison to the Same-ISA model baseline) in Appendix A.7.
Task-Specific Testing Datasets. The testing dataset for each ISA includes equal numbers of malware and benign samples. The benign samples are randomly selected from zlib, coreutils, and diffutils. Note that neither the benign nor malware samples are seen during the training of FlowMalTrans or the malware detection model.
Result Analysis. We first train LSTM on X86-64, and reuse it to test binaries in the other ISAs. The results in Table 3(a) show when the model trained on X86-64 is transferred to i386, ARM32, ARM64, MIPS32, PPC32 and M68K, it achieves AUC = $0.996$, $0.981$, $0.965$, $0.971$, $0.973$ and $0.931$, respectively. The high accuracies demonstrate the superior translation quality of FlowMalTrans.
Next, we analyze how FlowMalTrans preserves code semantics through translation. We visualize the token embeddings from different ISAs and find that tokens performing similar operations, regardless of their ISAs, have close embeddings. This demonstrates that FlowMalTrans effectively captures code semantics across ISAs. Further details are provided in Appendix A.8.
5.5 Model Comparison
We compare FlowMalTrans to the baselines and Same-ISA model (as described in Section 5.1).
Comparison with Baseline Methods. The first baseline is UnsuperBinTrans. As UnsuperBinTrans  focuses solely on two ISAs (X86-64 and ARM32), we train it on the same training datasets for the additional ISA pairs. We use it to translate binaries from the other ISAs to X86-64 and use LSTM to test the translated binaries. The results in Table 3(a) show that FlowMalTrans achieves better translation quality, leading to improved malware detection performance.
The second baseline is UniMap Wang et al. (2023a), which also aims to resolve the data scarcity issue in malware detection across ISAs. UniMap learns transferable knowledge for model reuse. It covers four ISAs, including X86-64, ARM32, MIPS32, and PPC32. To ensure a comprehensive comparison, we train it on the same datasets to cover the three additional ISAs (i386, ARM64 and M68K). Table 3(a) show that FlowMalTrans outperforms UniMap across all ISA pairs.
The third baseline is CrossIns2Vec Wang et al. (2024), which supports four ISAs (X86-64, ARM32, MIPS32, and PPC32). As it is not open-sourced, the comparison is limited to these four ISAs. The results show that FlowMalTrans achieves better performance than CrossIns2Vec.
The fourth baseline analyzes IR code. We assess whether a model trained on X86-64 IR code can be reused to test IR code in other ISAs. We use angr angr (2024) to lift binaries into IR. We train LSTM using the same task-specific training dataset in X86-64, and apply the model to test the same task-specific datasets in other ISAs. The AUC scores are $0.819$, $0.815$, $0.825$, $0.791$, $0.639$ and $0.476$ for i386, ARM32, ARM64, MIPS32, PPC32 and M68K, respectively. This indicates IR alone does not magically allow a model trained on one ISA to be reused for other ISAs. (more explanation is discussed in Appendix A.10).
Comparison with Same-ISA Model. The results of the Same-ISA model are shown in Table 3(b). When the LSTM model is trained and tested on the same ISA, it achieves AUC of $0.998$, $0.986$, $0.705$, $0.952$, $0.949$ and $0.790$ for i386, ARM32, ARM64, MIPS32, PPC32 and M68K, respectively. Comparing the results with our model in Table 3(a), we observe: (1) our model achieves performance close to Same-ISA model for i386, ARM32 and MIPS32; (2) our model significantly outperforms Same-ISA model for ARM64, PPC32 and M68K.
For i386, ARM32, and MIPS32, the results align with expectations: the Same-ISA model, trained and tested on the same ISA, outperforms FlowMalTrans, which is trained on X86-64 and tested on other ISAs (via translation). However, for ARM64, PPC32, and M68K, our model outperforms the Same-ISA model due to the insufficient malware samples used to train the Same-ISA model. For example, for M68K, only $80\%$ of $245$ malware samples are used for training. This highlights the importance of sufficient training data to achieve desirable performance. While including more data could enhance the model’s performance, collecting malware samples for low-resource ISAs can be challenging. Our approach—translating binaries to x86-64—addresses data scarcity effectively. In an extreme case, if only one binary in a given ISA is available, we can still detect whether it is malware using the X86-64-trained model.
5.6 Hyperparameter and Ablation Study
We conduct hyperparameter and ablation studies.
Normalizing Flows. We 1) explore different types of NFs, including scf (Scaling and Coupling Flow) and glow (Generative Flow), 2) vary the number of sequential flows, and 3) remove the flow adapter. The results show that 1) using scf with 3 sequential flows yields the best performance, and 2) when the flow adapter is removed, the performance degrades, highlighting the crucial role of normalizing flows.
Normalization Rules. We conduct ablation study to evaluate the impacts of normalization rules (Section 4.1). The results show that: (1) When all rules are applied, we achieve the best performance. (2) When a subset of rules is applied, the AUC values are reduced, indicating each rule mitigates the OOV issue and positively impacts translation quality.
More details are presented in Appendix A.9.
6 Conclusion and Future Work
We proposed FlowMalTrans, a flow-adapter-based unsupervised binary code translation model. FlowMalTrans incorporates normalizing flows (NFs) to model the basic block-level representations, enhancing its code translation capability. We train FlowMalTrans to translate binaries from other ISAs to X86-64, and reuse a malware detection model trained on X86-64 to test the translated code. Our approach effectively reduces the burden of data collection and achieves better malware detection across ISAs than prior state of the arts.
Large language Models (LLMs) show extraordinary capability in coding tasks and are popular utilized Xu et al. (2025); Ma et al. (2024). We leave using LLM for code translation as future work.
Limitations
Generalizability. Our evaluation encompasses malware detection across seven distinct ISAs, demonstrating the broad applicability of our method. However, our experimental validation is constrained to Linux-based malware samples. The transferability of our binary code translation framework to alternative operating systems—including Windows, iOS, and Android environments—remains an open question that warrants dedicated empirical investigation. Additionally, while our experiments focus on translating from various ISAs to X86-64, exploring ARM32 as an alternative target architecture presents an interesting direction, given its substantial malware dataset availability.
Beyond malware detection, our translation framework shows promise for broader binary analysis applications. The underlying code transformation capabilities could be leveraged for tasks such as vulnerability discovery and code similarity analysis. The strong performance demonstrated in malware detection validates the core technical approach, establishing a foundation for future research directions. Expanding this methodology to additional security analysis tasks represents a natural progression that could unlock significant research opportunities in cross-ISA binary analysis.
Malware Packing and Obfuscation. The practice of malware packing entails hiding malicious code inside files that appear benign. Our research excludes malware packing from consideration, as it lies beyond our analytical scope. Our investigation centers on examining how well malware detection models transfer across different instruction set architectures (ISAs) through binary translation. The malware specimens employed in our detection experiments are uncompressed, enabling their examination and reverse engineering using analytical tools such as IDA Pro IDA (2023).
Additionally, when dealing with compressed malware, one approach involves utilizing sophisticated decompression utilities, including PEiD PEiD (2008) and OllyDbg OllyDbg (2000), to initially decompress the malware before analyzing the resulting content.
Code obfuscation methods present significant obstacles for malware identification systems. The malware specimens utilized in our research originated from VirusShare.com VirusShare (2020), a collection platform that gathers malware encountered in real-world environments. It is commonly understood that such malware frequently implements obfuscation strategies to circumvent detection systems. Nevertheless, we do not possess definitive information about the particular obfuscation methods employed in each malware sample, which complicates evaluating resistance against specific obfuscation approaches.
Our research primarily tackles the problem of limited data availability in resource-constrained ISAs by converting binaries from these architectures to a well-resourced ISA through FlowMalTrans.
Subsequent research could methodically investigate how obfuscation techniques affect detection capabilities. A significant obstacle is the lack of a reliable, high-quality dataset that correlates malware samples with their corresponding obfuscation methods. Addressing this deficiency will be a priority in our upcoming investigations.
Ethical Considerations
To train and test our translation model FlowMalTrans, we first collect various open-source programs and compile them for different ISAs using cross compilers. Given the wide availability of open-source programs, this requires minimal effort. Moreover, we strictly adhere to the licensing agreements and intellectual property rights associated with each program when collecting them and building the training and testing datasets.
For evaluation within the security domain, we utilized malicious software specimens obtained from the VirusShare.com platform VirusShare (2020), a well-established digital repository that serves the cybersecurity research community. We acknowledge the substantial ethical responsibilities accompanying the use of such security-sensitive materials. Our research protocols emphasized legitimate scientific inquiry, prevention of inadvertent threat proliferation, and careful attention to confidentiality and regulatory compliance.
In the interest of scientific reproducibility and community advancement, we intend to distribute our developed model architecture, implementation code, and experimental datasets through a public repository under GPL licensing terms. The compiled open-source program collections will be made fully accessible in their original form. Regarding security-critical materials, we will exclusively share cryptographic hash identifiers and file references from VirusShare.com, enabling verification and retrieval by qualified researchers while avoiding direct transmission of potentially dangerous executable content.
Appendix A Appendix
A.1 Model Training Details
Model Pretraining. Pretraining is a key ingredient of unsupervised neural machine translation. Studies have shown that the pre-trained cross-lingual word embeddings which are used to initialize the lookup table, have a significant impact on the performance of an unsupervised machine translation model Conneau et al. (2020); Conneau and Lample (2019). We adopt this and pre-train both the encoders and decoders of FlowMalTrans to bootstrap the iterative process of our binary translation model. We employ causal language modeling (CLM) and masked language modeling (MLM) objectives to train the encoder and decoder.
(1) The CLM objective trains the model to predict a token $e_{t}$, given the previous (t1)(t-1) tokens in a basic block $P(e_{t}|e_{1},...,e_{t-1},\theta)$. (2) For MLM, we randomly sample $15\%$ of the tokens within the input block and replace them with [MASK] $80\%$ of the time, with a random token $10\%$ of the time, or leave them unchanged $10\%$ of the time.
The first and last token of an input basic block is a special token [/s], which marks the start and end of a basic block. We add position embedding and architecture embedding to token embedding, and use this combined vector as the input to the bi-directional Transformer network. Position embeddings represent different positions in a basic block, while architecture embeddings specify the architecture of a basic block. Both position and architecture embeddings are trained along with the token embeddings and help dynamically adjust the token embeddings based on their locations.
Denoising Auto-Encoding (DAE). The DAE reconstructs a basic block from its noised version, as depicted in the process 1 and 2 in Figure 2. Given the input source block, $B_{src}$, we introduce random noise into it (e.g., altering the token order by making random swaps of tokens), resulting in the noised version, $B^{{}^{\prime}}_{src}$. Then, $B^{{}^{\prime}}_{src}$ is fed into the target encoder, whose output is analyzed by the source decoder. The training aims to optimize both the target encoder and source decoder to effectively recover $B_{src}$. Through this, the model can better accommodate the inherent token order divergences. Similarly, the source encoder and source decoder are optimized when the input is a target basic block, $B_{tgt}$. The training procedure of DAE involves only a single ISA at each time, without considering the final goal of translating across ISAs nor the latent code transformation.
Back Translation (BT). We adopt the back-translation approach Feldman and Coto-Solano (2020) to train our model in a translation setting, as shown in Figure 2. As an example, given an input basic block in one ISA $N$, we use the model in the inference mode to translate it to the other ISA $M$ (i.e. applying the encoder of $N$ and the decoder of $M$). This way, we obtain a pseudo-parallel basic block pair. We then train the model to predict the original basic block (i.e. applying the encoder of $M$ and the decoder of $N$) from this synthetic translation. As training progresses and the model improves, it will produce better synthetic basic block pairs through backtranslation, which will serve to further improve the model in the subsequent iterations. In BT, the latent code transformation is performed twice and trained along with the encoders/decoders; taking BT for the source ISA as an example: first in the source-to-target direction, then in the target-to-source direction as shown in the steps of 3 and 4 in Figure 2.
Table 4: Comparison of original and normalized assembly code.
call _gpgrt_log_info call <FUNC>
sbb al 0 sbb al <VALUE>
mov esi 0ACh+_bss_start mov esi <HEX>+<TAG>
lea rdi str_LogWithPid lea rdi <STR>
sub rdx buffer sub rdx <TAG>
(a) X86-64
add esp 0Ch add esp <HEX>
call _dcgettext call <FUNC>
lea eax [ebx-5D40h] lea eax [ebx-<HEX>]
jmp short loc_37C3 jmp short <LOC>
mov eax [150h+domainname] mov eax [<HEX>+<TAG>]
(b) i386
ADD R12 R12 0x1B000 ADD R12 R12 <HEX>
LDR PC memcpy-0x2B7C LDR PC <FUNC>-<HEX>
BEQ.W loc_109B4 BEQ.W <LOC>
BL gz_uncompress BL <FUNC>
CMP R2 0 CMP R2 <VALUE>
(c) ARM32
LDR X0 [SP #0xC0+stream_68] LDR X0 [SP <HEX>+<TAG>]
TBZ W0 #0 loc_AF38 TBZ W0 <HEX> <LOC>
ADRL X1 str_ErrorInitia ADRL X1 <STR>
B _gmon_start_ B <FUNC>
MOV X19 #0 MOV X19 <HEX>
(d) ARM64
lw \$fp 0x40+var_20 lw \$fp <HEX>+<VAR>
bal usage bal <FUNC>
beqz \$v0 loc_1358 beqz \$v0 <LOC>
move \$a2 \$s3+1 move \$a2 \$s3+<VALUE>
sw \$s0 0x40+path(\$sp) sw \$s0 <HEX>+<TAG>(\$sp)
(e) MIPS32
addi r3 r4 8 addi r3 r4 <VALUE>
cmpwi r2 8+var_12 cmpwi <VALUE>+<VAR>
beq _str_branch_ beq <TAG>
stw r6 r2 6+0x1C5 stw r6 r2 <VALUE>+<HEX>
lwz r3 0x3H+str_56 lwz r3 <HEX>+<STR>
(f) PPC32
move.l #\$DB3EF d0 move.l <HEX> d0
sub.w a2 d4-stream_23 sub.w a2 d4-<TAG>
tst.b loc_5H tst.b <LOC>
cmp.l #\$352A #23+mem_name cmp.l <HEX> <VALUE>+<TAG>
divs.w #16 d2 divs.w <VALUE> d2
(g) M68K
A.2 Examples of Normalized Assembly Code
Table 4 shows some randomly selected examples for instruction normalization. For each sub-table, the code on the left side represents the original version, while the code on the right side shows the normalized version.
A.3 Dataset Adequacy
To train FlowMalTrans, we create a training dataset for each ISA: $2,789,119$ blocks for X86-64, $2,803,557$ blocks for i386, $7,413,083$ blocks for ARM64, $5,812,795$ blocks for ARM32, $4,813,685$ blocks for MIPS32, $3,964,237$ blocks for PPC32, and $5,463,395$ blocks for M68K. It should be noted that the training of FlowMalTransis unsupervised.
In NLP, it is widely recognized that a comprehensive dataset, which ensures that the vocabulary covers a wide range of words, is crucial for training effective code translation models. We assess the adequacy of our mono-architecture datasets. Specifically, we study the vocabulary growth as we incrementally include programs. Our findings indicate that while the vocabulary size initially increases with the inclusion of more programs, it eventually stabilizes. Take x86-64 as an example, including openssl-1.1.1p results in a vocabulary size of $23,029$. The size increases to $36,770$ (a $60\%$ growth) when binutils-2.34 is added, and then increases to $39,499$ (a $7.4\%$ growth) and $39,892$ (a $0.9\%$ growth) when findutils-4.8.0, libpgp-error-1.45 are included, respectively. The growth trend is similar for other ISAs. It shows that the vocabulary barely grows in the end when more programs are added. According to the vocabulary growth trend and the high performance achieved, our mono-arch datasets are adequate to cover instructions and enable effective code translation.
Note that the datasets used for training FlowMalTrans have no overlap with (1) the dataset used for testing the translation capability of FlowMalTrans and (2) the testing dataset used in the malware detection task. The details of these datasets are introduced in the following sections.
A.4 Model Parameters of FlowMalTrans
The encoders and decoders of FlowMalTrans are implemented using the Transformer model. Table 5 shows the details of the model parameters.
Table 5: Parameter Details of FlowMalTrans.
Parameter Value Description
Emb. Dimension 32/64/128 Embedding layer size for tokens
Hidden Dimension 4* Emb. Dimension Transformer FFN hidden dimension
Num. Layers 4 Number of transformer layers
Num. Heads 4 Number of attention heads per layer
Regu. Dropout 0.1 Dropout rate for regularization
Attn. Dropout 0.1 Dropout rate in attention layers
Batch Size 256 Number of sentences per batch
Max. Length 512 Maximum length of one sentence after BPE
Optimizer Adam Adam optimizer with sqrt decay
Clip Grad. Norm 5 Maximum gradient norm for clipping
Act. Function ReLU Use ReLU for activation
Pooling Mean Use mean pooling for sentence embeddings
Accumulate Grad. 8 Accumulate gradients over N iterations
A.5 Model Parameters of Malware Detection LSTM Model
We use the LSTM model to detect malware. Table 6 shows the parameters details of the LSTM model.
Table 6: Parameter Details of the malware detection LSTM model.
Parameter Value Description
Emb. Dimension 32/64/128 Input feature dimension for sequence embedding
Num. of Layers 3 Number of stacked LSTM layers in the network
Hidden Units 16 Number of hidden units in each LSTM layer
Output Units 1 Dimension of the output layer
Batch Size 36 Number of samples processed in one batch
Optimizer Adam Adaptive optimization algorithm with momentum
Loss Function BCEWithLogitsLoss Binary cross-entropy with logits
Pooling Max Maximum value across temporal dimension
Table 7: Parameter Details of the malware detection CNN model.
Parameter Value Description
Emb. Dimension 64 Input feature dimension for sequence embedding
Conv. Layers 2 Number of convolutional layers in the CNN network
Conv. Kernel 3 Kernel size of the convolutional layers
Output Units 1 Dimension of the output layer
Batch Size 64 Number of samples processed in one batch
Optimizer Adam Adam optimizer with a learning rate of 0.001
Loss Function BCEWithLogitsLoss Binary cross-entropy with logits
Pooling Max Maximum pooling to reduce spatial dimensions
A.6 Malware Detection Using a CNN model
We use a 1-dimensional Convolutional Neural Network (CNN) model to detect malware. The parameters of the CNN model are presented in Table 7.
We use the same task-specific training and testing datasets described in Section 5.4. Moreover, we compare our results against four baseline methods and the Same-ISA model. Table 8 presents the malware detection results using the CNN model.
Table 8: Malware detection results using CNN. In Table (a), we compare the detection performance by translating malware using FlowMalTrans, UnsuperBinTransṪhen we evaluate it against UniMap and the IR-based and Same-ISA model. In Table (b), we give the results of the Same-ISA model, which is trained and tested on the same ISA.
(a) FlowMalTrans vs. Four baselines.
ISA Pair FlowMalTrans UnsuperBinTrans UniMap CrossIns2Vec IR-based Model
(src $\rightarrow$ tgt) (Our Work) (Baseline 1) (Baseline 2) (Baseline 3) (Baseline 4)
i386 $\rightarrow$ X86-64 0.993 $0.745$ $0.953$ - $0.823$
ARM32 $\rightarrow$ X86-64 0.984 $0.826$ $0.953$ $0.976$ $0.842$
ARM64 $\rightarrow$ X86-64 0.973 $0.641$ $0.910$ - $0.818$
MIPS32 $\rightarrow$ X86-64 0.974 $0.762$ $0.933$ $0.964$ $0.774$
PPC32 $\rightarrow$ X86-64 0.972 $0.736$ $0.919$ $0.967$ $0.642$
M68K $\rightarrow$ X86-64 0.931 $0.696$ $0.893$ - $0.483$
(b) The Same-ISA model.
Train & Test on Optimal
the same ISA Model
i386 $0.997$
ARM32 $0.988$
ARM64 $0.714$
MIPS32 $0.956$
PPC32 $0.853$
M68K $0.780$
When the CNN model trained on X86-64 is transferred to i386, ARM32, ARM64, MIPS32, PPC32 and M68K, it achieves AUC values of $0.993$, $0.984$, $0.973$, $0.952$, $0.942$ and $0.937$, respectively. These high accuracies highlight the superior translation quality of FlowMalTrans, outperforming both UnsuperBinTrans and the IR-based model. Compared to the Same-ISA model, we have the following observation. (1) First, our model achieves performance close to the Same-ISA model when testing malware on i386, ARM32, and MIPS32. This outcome is expected, as the Same-ISA model is trained and tested on the same ISA, while our model is trained on X86-64 and tested on other ISAs through translation. (2) Second, our model significantly outperforms the Same-ISA model on ARM64, PPC, and M68K. This is due to the limited malware samples available for training the Same-ISA model on the two ISAs, highlighting the value of our approach. By reusing a model trained on a high-resource ISA, we enable robust detection for low-resource ISAs that would otherwise face significant challenges.
A.7 Malware Detection Using All Malware Samples for Testing
We train the LSTM model exclusively on X86-64, and reuse the trained model to test binaries in other ISAs, including i386, ARM32, ARM64, MIPS32, PPC32 and M68K. It is important to note the key difference between the experiment described in this Appendix and that in Section 5.4. Here, we use all malware samples from i386, ARM32, ARM64, MIPS32, PPC32 and M68K for testing. In contrast, in Section 5.4, only 20% of the malware samples from these ISAs are used for testing, as the remaining 80% are reserved for training.
Table 9: Malware detection results. We use all the malware samples in i386, ARM32, ARM64, MIPS32, PPC32 and M68K for testing. We compare the detection performance by translating malware using FlowMalTrans and UnsuperBinTrans, and evaluate it against the IR-based model.
ISA Pair FlowMalTrans UnsuperBinTrans UniMap CrossIns2Vec IR-based Model
(src $\rightarrow$ tgt) (Our Work) (Baseline 1) (Baseline 2) (Baseline 3) (Baseline 4)
i386 $\rightarrow$ X86-64 0.997 $0.728$ $0.949$ - $0.821$
ARM32 $\rightarrow$ X86-64 0.982 $0.821$ $0.951$ $0.975$ $0.813$
ARM64 $\rightarrow$ X86-64 0.963 $0.641$ $0.913$ - $0.822$
MIPS32 $\rightarrow$ X86-64 0.974 $0.731$ $0.931$ $0.968$ $0.796$
PPC32 $\rightarrow$ X86-64 0.978 $0.732$ $0.916$ $0.973$ $0.641$
M68K $\rightarrow$ X86-64 0.940 $0.682$ $0.894$ - $0.480$
Result Analysis. We first train LSTM on X86-64, and reuse the model to test binaries in the other ISAs. The results are shown in Table 9. We can see that when the model trained on X86-64 is transferred to i386, ARM32, ARM64, MIPS32, PPC32 and M68K, it achieves AUC = $0.997$, $0.982$, $0.963$, $0.974$, $0.978$ and $0.940$ respectively. The high accuracies demonstrate the superior translation quality of FlowMalTrans.
A.8 Semantic Transfer Analysis
Figure 3: Visualization of opcode embeddings.
We analyze how FlowMalTrans is able to preserve the code semantics through translation. Specifically, we visualize the embeddings of opcode tokens from different ISAs. We take the X86-64 and ARM32 pair as an example. Opcodes, which determine the operation to be performed, capture more semantics compared to operands; we thus focus on opcodes for this demonstration. We extract the embeddings of $138$ X86-64 opcodes and $247$ ARM32 opcodes from FlowMalTrans, and visualize them using t-SNE, as shown in Figure 3. Four categories of opcodes are selected for demonstration. We can see that opcodes performing similar operations, regardless of their ISAs, are close together. Thus, FlowMalTrans can successfully capture semantic relationships of opcodes across ISAs and preserve code semantics.
A.9 Hyperparameter and Ablation Study
Table 10: Impacts of different types of normalizing flows on the malware detection task.
Type i386 ARM32 ARM64 MIPS32 PPC32 M68K
3-scf 0.996 $0.981$ 0.965 0.971 0.973 0.931
3-glow $0.994$ 0.983 $0.924$ $0.823$ $0.873$ $0.873$
5-scf $0.957$ $0.875$ $0.817$ $0.910$ $0.832$ $0.824$
5-glow $0.919$ $0.891$ $0.843$ $0.804$ $0.821$ $0.760$
None $0.854$ $0.871$ $0.784$ $0.769$ $0.811$ $0.752$
Normalizing Flows. We conduct an experiment to analyze the impact of normalizing flows on translation capability based on the following aspects: (1) exploring different flow-adapter architectures, such as Glow Diederik P. Kingma (2014); (2) varying the number of sequential flows to build the source/target flow; and (3) removing the flow-adapter architecture in FlowMalTrans.
The results are shown in Table 10. We use two types of normalizing flows: scf (Scaling and Coupling Flow) and glow (Generative Flow), each configured with different numbers of sequential flows (e.g., 3 and 5). For instance, 5-glow represents the scenario where five consecutive glow flows are utilized to construct the source/target flow. The last row, None, represents the scenario in which no normalizing flows are incorporated into the model FlowMalTrans.
We can observe that: (1) Using 3-scf to construct the normalizing flows yields the best performance across all ISAs except for ARM32. However, the result for ARM32 remains satisfactory, being close to the highest value. (2) When the flow-adapter architecture is removed from FlowMalTrans, the performance degrades, highlighting the crucial role of normalizing flows in enhancing translation capability.
Normalization Rules. Each instruction in the datasets is normalized by applying the three rules (R1, R2, and R3) discussed in Section 4.1. Normalization is a vital step in our approach. In this experiment, we conduct ablation study by removing certain rules and evaluating their influence on malware detection. We consider these cases:
  • (C1): Applying all rules to the data.
  • (C2): Removing R1, applying R2 and R3.
  • (C3): Removing R2, applying R1 and R3.
  • (C4): Removing R3, applying R1 and R2.
  • (C5): Not applying any rules to the data.
Table 11: AUC when varying normalizing strategies.
Case # i386 ARM32 ARM64 MIPS32 PPC32 M68K
Case 1 0.996 0.981 0.965 0.971 0.973 0.931
Case 2 $0.862$ $0.521$ $0.624$ $0.623$ $0.617$ $0.638$
Case 3 $0.883$ $0.575$ $0.617$ $0.570$ $0.503$ $0.642$
Case 4 $0.827$ $0.533$ $0.613$ $0.422$ $0.451$ $0.546$
Case 5 $0.793$ $0.453$ $0.523$ $0.422$ $0.473$ $0.471$
Table 11 shows the results. We can observe that: (1) When all normalization rules are applied (C1), we achieve the best performance. (2) When a subset of normalization rules is applied (C2-4), the AUC values are lower than in C1, indicating that each normalization rule mitigates the OOV issue and has an impact on translation quality, thereby influencing malware detection performance. (3) When no normalization rule is applied (C5), the results are the lowest.
Embedding Dimension. We evaluate the impacts of the embedding dimension. We test different dimension sizes, including $32$, $64$, and $128$, to train FlowMalTrans. We then apply FlowMalTrans to translate binaries from the source ISA to X86-64 for malware detection. The results are shown in Table 12. We observe that when the dimension is set to $64$, the AUC values are higher compared to other dimensions. Moreover, as the dimension increases, the training time also increases. We thus choose a dimension of $64$, considering both the translation quality and training efficiency.
Table 12: AUC when varying dimension size.
ISA Pair Dimension: Dimension: Dimension:
(src $\rightarrow$ tgt) $32$ $64$ $128$
[…]
A.10 Discussion on IR-based model
Although IR can abstract away many architectural differences among ISAs, significant variations persist across IR code from different ISAs. Given two binary executables from different ISAs, compiled from the same source code, their corresponding IR representations can still appear quite different. This issue is discussed in Section 4.3.1 of the UniMap paper. Consequently, existing approaches that use IR for cross-ISA binary analysis often require additional advanced techniques, such as fuzzing or re-optimization, to bridge these differences. Similarly, in our evaluation, the IR-based baseline, which trains a model on x86 IR code and reuses it for testing on IR code from other ISAs, performs poorly due to these variations.
Decompilation, which converts binary code into readable C-like code, is inherently a lossy process with limited accuracy (e.g., complex constructs such as optimized loops, switch statements, and inlined functions are often not recovered accurately). Moreover, while IDA Pro’s decompiler is among the best available, it supports only a limited set of ISAs, restricting its generalization to only those architectures. In contrast, our binary code translation approach does not suffer from these limitations. Thus, decompilation is an unsuitable solution for our needs.
A.11 Translation Demonstration
Table 13 and Table 14 shows some randomly selected examples. In each example, (1) the source ISA is the original basic block in the source ISA, which could be i386, ARM32, ARM64, MIPS32, PPC32 or M68K; (2) the target ISA is the basic block in the target ISA, X86-64, that is similar to the original basic block in the source ISA; and (3) the translated ISA is the X86-64 basic block translated from the original basic block in the source ISA by our model FlowMalTrans.
By comparing the translated X86-64 block with the target X86-64 block, we observe that FlowMalTrans (1) accurately predicts almost all opcodes, and (2) while a few operands are predicted incorrectly, these errors are reasonable. Note that an instruction consists of an opcode (which specifies the operation) and zero or more operands (which specifies registers, memory locations, or literal data). Thus, opcodes, which determine the operation to be performed, capture more semantic information compared to operands. On the other hand, different registers or memory locations can store data while preserving code functionality, which reduces the significance of operands.
[…]
Table 13: Examples for code translation.
Source i386
sub esp <VALUE> add edx [s+<HEX>] mov ebx esi push len push edx
push [esp+<HEX>+dictionary] call <FUNC> add esp <HEX> jmp <LOC>
1 Translated X86-64
sub rsp <VALUE> add rdx [s+<HEX>] mov rbx rbp push len
mov r8 [rsp+<HEX>+datalen] mov rcx len call <FUNC> jmp <ADDR>
Target X86-64
sub rsp <VALUE> add rbp [state+<HEX>] mov eax <HEX> push <TAG>
mov [state+<HEX>] rdi mov rdi state call <FUNC> jmp short <TAG>
Source i386
mov edx [edi+<VALUE>] mov edp <VALUE> cmp edx <HEX> lea ecx [edx-<HEX>]
setnz al and ecx <HEX> lea ecx [edx-<HEX>]
i386 2 Translated X86-64
mov rdx [abfd+<VALUE>] mov ebp <VALUE> cmp rdx <HEX> lea rcx [r1+rax+<VALUE>]
setnz al and ecx <HEX> lea rcx [r1+rax+<VALUE>]
Target X86-64
mov rdx rsp mov rsp <HEX> cmp rax rdx lea r8 <TAG>
setnz a1 and ecx <TAG> lea rcx [rax+<HEX>]
Source i386
sub esp <VALUE> lea eax [esi+<VALUE>] push eax push [esp+<HEX>+buf]
call <FUNC> mov edi eax mov c [esp+<HEX>+<TAG>]
3 Translated X86-64
sub rsp <VALUE> lea rax <TAG>+<HEX> mov cs:<TAG> rax mov rax [abfd+<VALUE>]
call <FUNC> mov rsi rsp mov rdi abfd call <FUNC> add rsp <VALUE>
Target X86-64
sub rsp <HEX> lea eax <HEX> mov rax rs:<HEX> mov [rsp+<HEX>+<VAR>]
call <FUNC> mov rcx rsp mov rsi <VALUE> call <FUNC> add eax <HEX>
Source ARM32
XOR R0 R2 MOV R1 R1 <VALUE> SUBS R3 R0 R3
LDP R6 R6 have BNE <TAG> CMP copy R7
1 Translated X86-64
xor edi edi mov rcx rdx and esi <VALUE> sub <TAG> <VALUE>
mov [rax+rdx-<VALUE>] cl mov ecx <VALUE>
Target X86-64
xor edi edi mov [rax+rcx-<VALUE>] di and edi <TAG> sub init curr
mov [s+<HEX>] rdx mov rsp <VALUE>
Source ARM32
LDR R2 [R3] MOVS R0 <VALUE> BNE <LOC>
LDR R0 [R2+<TAG>] MOV R1 <TAG> BL <FUNC>
ARM32 2 Translated X86-64
mov rsi [rsp+<HEX>+p] movsxd rdi eax
lea rcx [rsp+<HEX>+id] call <FUNC>
Target X86-64
mov rdi in movsxd rdi file
lea rax [rsi+<VALUE>] call <FUNC>
Source ARM32
MOV R0 strm MOVS R1 <VALUE> BL <FUNC> ADDS err <VALUE>
MOVS R2 <TAG> LDR R1 <HEX> BNE <ADDR>
3 Translated X86-64
mov edi ebx lea rcx <TAG> movsxd rax ds:<TAG> call <FUNC>
add rax rcx lea rdx <TAG> jmp <ADDR>
Target X86-64
mov edi ebx lea rsi <HEX> movzx rax <TAG> call <FUNC>
add eax ecx lea rbp <LOC> jmp <TAG>
Source ARM64
MOV W2 <VALUE> MOV W0 W2 LDP X2 <VALUE> [SP+<HEX>+<TAG>]
LDR X2 [SP+<HEX>+<TAG>] LDP X2 X3 [SP+<HEX>+<TAG>]
1 Translated X86-64
mov qword ptr [rax] <VALUE> mov rax [rbp+p]
mov rdi rax call <FUNC> test eax eax jnz <ADDR>
Target X86-64
mov qword ptr [rcx+<HEX>] <VALUE> mov [rdx+<HEX>] <VALUE>
mov [rdx+<HEX>] rcx call <FUNC> test rax rax jnz <TAG>
Source ARM64
ADD W2 W2 <VALUE> ADD X3 X0 X3 SUB W2 W2 <VALUE>
MOV W0 <VALUE> STRB W0 [X1+<HEX>] MOV X0 X2 <VALUE>
ARM64 2 Translated X86-64
mov r1 q mov rsi p movzx ri byte ptr [p]
sub r1 <VALUE> add p <HEX> mov rax [p] jmp short <LOC>
Target X86-64
mov r14 rsp mov rsip movzx rax <HEX>
sub r10 <VALUE> add q <HEX> mov r11 <VALUE> jmp <TAG>
Source ARM64
LDR W5 <VALUE> LDR X9 <HEX> SUB W0 W5 W0 ADD W1 W0 <VALUE> LDP W1 <VALUE>
SUB W0 W5 W0 ADD W1 W0 <VALUE> LDP W1 <VALUE> LDRB W2 <TAG>
3 Translated X86-64
mov rdx <HEX> mov [rsp+<HEX>+n] rdx movzx edx [r1+<HEX>]
mov esi edx mov r9 [r1+<HEX>] lea r1 [rdx+<TAG>]
Target X86-64
mov rdx [rbp+mode] mov ecx [rbp+fd] movzx rax [rbp+path] mov esi ecx
mov rdi rax mov [rbp+gz] rax lea r1 [rax+<VALUE>]
Table 14: Examples for code translation.
Source MIPS32
li \$t9 <VALUE> lw \$ra <HEX>+<VAR> (\$sp)
addiu \$t9 <FUNC> b <FUNC> addiu \$sp <HEX>
1 Translated X86-64
mov r8 <VALUE> mov rbx rsp+<HEX>+<TAG>
lea r10 r8+<ADDR> call <FUNC> add rbp <VALUE>
Target X86-64
mov r10 <VALUE> mov rax rsp+<HEX>+<VAR>
lea r10 r10+<TAG> jmp <FUNC> add rsp <HEX>
Source MIPS32
li \$a1 <VALUE> addiu \$a1 <STR> move \$a0 \$s0
la \$t9 <FUNC> jalr \$t9 lw \$gp <HEX>+<VAR> (\$sp)
MIPS32 2 Translated X86-64
mov rsi <HEX> lea rbp rbp+<ADDR> mov rdi rax
mov r11 <FUNC> call <FUNC> mov rap rsp
Target X86-64
mov rsi <VALUE> lea rsi rsi+<STR> mov rdi, rbx
mov r10 <ADDR> call <FUNC> mov rbp rsp
Source MIPS32
li \$a2 <VALUE> li \$a1 <VALUE> addiu \$a1 <STR> move \$a0 \$zero la \$t9 <FUNC>
jalr \$t9 lw \$gp <HEX+<VAR> (\$sp) move \$a1 \$s1 move \$a0 \$v0
3 Translated X86-64
mov rbx <HEX> mov rsi <ADDR> lea rsi rsi+<TAG> xor rsi rsi
mov r10 r9 call <FUNC> mov rbp rap mov rsi r14+<TAG>+<HEX> mov rdi rax
Reference X86-64
mov rdx <VALUE> mov rsi <VALUE> lea rsi rsi+<STR> xor rdi rdi
mov r10 r11 call <FUNC> mov rbp rsp+<HEX>+<VAR> mov rsi r14 mov rdi rax
Source PPC32
lis r13 r2@ha addi r3 r3 at1@l lis r9 <VAR>+<HEX>@ha lwz r1 <TAG>+<LOC>@l(r9)
add r10 r7 r8 add r10 r6 r10 lwz r3 <HEX>(r1) bl <FUNC>
1 Translated X86-64
mov rbi rax movsz rbp [<HEX>+<VALUE>] mov rdx rax lea [rax r1+<TAG>]
mov rdx [rsi+<LOC>] call <FUNC> test rax rax jz <TAG>
Target X86-64
mov rbi <LOC> mov rbx <VALUE> lea [rdi+<LOC>]
mov rdx [rbx+<VALUE>] call <FUNC> test rdx rdx jnz <LOC>
Source PPC32
mr r3 r7 li r3 <VAR> lis <FUNC>
li r6 <TAG> mr r5 r4 stwu r31 -4(r1) b <LOC>
PPC32 2 Translated X86-64
movsx r2 r13 mov rbi <VAR> call <FUNC>
mov rbx <TAG> mov rax rdi push rbx jmp <ADDR>
Target X86-64
mov r12 rbi movsz rsi mov rbx <LOC> lea <HEX>+<TAG>
mov r12 r7 mov rbx (<VALUE>+<HEX>) mov r8 <HEX> jmp <HEX>
Source PPC32
extsh r2 r13 li r9 <VAR> bl <FUNC> lis r5 <TAG>@ha
addi r5 r5 <TAG>@l mr r6 r3 stwu r5 -4(r1) b <ADDR>
3 Translated X86-64
mov rbx <LOC>+<HEX> movsx rbx rax mov eax [ecx+<HEX>]
mov ebi [ecx+<HEX>] mov r5 [r2+<VALUE>] lea r7 [rax+<HEX>]
Reference X86-64
mov rax [<VALUE>+<HEX>] mov eax <VAR> mov rbp [rsi+<VALUE>] mov eax ebx
mov rsi rdx mov rax (rax+<VAR>) l rbi [r3+<VAR>]
Source M68K
move.l at2 a0 move.l <VALUE>+<TAG> a7 move.l a1 d2 add.l a2 d3
add.l a3 d2 move.l <HEX>(a7) a0 jsr <FUNC> tst.l d0 bne <LOC>
1 Translated X86-64
mov rsi at3 mov rsp [<HEX>+<TAG>] mov rax rex lea [r4 r10+r3]
mov rbx [rbp+<VALUE>] call <FUNC> test rax rax jz <LOC>
Target X86-64
mov rsi <TAG> mov rbp <VALUE> lea [rdx+<HEX>]
mov rdi [rax+<VALUE>] call <FUNC> test rax rax jnz <TAG>
Source M68K
move.l d3 d2 move.l <HEX> a0 lea <FUNC> a1
move.b <TAG> d4 ext.l d4 move.l d1 d0 move.l a6 a7 b <LOC>
M68K 2 Translated X86-64
mov r12 rbx mov rdi <HEX> lea <FUNC> movsx rax <TAG>
mov rbx rdx push rbp jmp short <LOC>
Target X86-64
mov r12 rbp mov rsp movsx rcx <LOC> lea <TAG>
mov r2 r4 mov rbp (<TAG>+<HEX>) mov r11 <VALUE> jmp <ADDR>
Source M68K
move.l <VALUE>+<VAR> d0 move.l d2 d1 move.l <VALUE>(d1) d2 move.l d1 a0
move.s d3 <VALUE> move.l <HEX>(a1) d3 lea <TAG>(a2) a1
3 Translated X86-64
mov rax <VAR>+<HEX> mov rcx rbx mov ebx [ecx+<VALUE>]
mov ebi ecx mov r9 [r1+<HEX>] lea r1 [rdx+<TAG>]
Reference X86-64
mov rax [<VALUE>+<HEX>] mov eax <VAR> mov rbp [rsi+<VALUE>] mov eax ebx
mov rsi rbx mov rax (rax+<TAG>) lea rsi [r2+<VAR>]
A.12 Uniqueness of Our Work
We propose an entirely different approach compared to the prior works Wang et al. (2023b, 2024), such as UniMap Wang et al. (2023b), which also aims to reuse models across ISAs in order to resolve the data scarcity issue in low-resource ISAs.
A key limitation of UniMap is its reliance on a linear transformation to map instruction embeddings between ISAs. This approach assumes that the embedding spaces of different ISAs exhibit structural similarity (i.e., are isomorphic). However, this assumption may not hold, particularly for ISAs with significantly different vocabulary sizes, making it infeasible to find a suitable linear transformation. In contrast, FlowMalTrans does not rely on such an assumption, making it more broadly applicable across diverse ISAs.
While UniMap learns cross-architecture instruction embeddings, our method focuses on translating binary code across ISAs. By translating code to a high-resource ISA, our approach offers several advantages. First, it allows the direct application of existing downstream models—trained on the high-resource ISA—to other ISAs through testing the translated code. In contrast, the works in Wang et al. (2023b, 2024) require retraining the model using cross-architecture instruction embeddings. Furthermore, translating code from one ISA to another assists human analysts in understanding code from unfamiliar ISAs, supporting broader applications in code comprehension.
InnerEye Zuo et al. (2018) applies neural machine translation (NMT) techniques for binary code similarity comparison but does not perform binary code translation across ISAs. It uses two encoders from NMT models, where each encoder generates an embedding for a piece of binary code of a given pairs, and measures similarity based on embedding distance. In contrast, our approach focuses on translating binary code across ISAs.
Intermediate Representation (IR) is an architecture-agnostic representation that abstracts away various architectural differences among ISAs. However, despite this abstraction, significant variations still exist across IR code from different ISAs. Consequently, even when binaries from different ISAs—compiled from the same source code—are converted into a common IR, the resulting IR code often differs significantly, as discussed in Ahmad and Luo (2023). We compare malware detection performance across ISAs by translating binaries using FlowMalTrans against an approach that analyzes the IR code of binaries across ISAs. The results show our model enables superior cross-ISA malware detection (see the evaluation). An interesting direction is exploring translation based on IR code, which we leave for future work.